OpenStudy (anonymous):

Please help! :) In basketball, a field goal is a basket scored on any shot other than a free throw. It may be worth two or three points, depending on the shooter’s distance from the basket. In a recent game, Max’s team scored T three-point field goals and W two-point field goals. a. How many points did the team score from field goals? b. If the team made 7 three-point field goals and earned a total of 67 points in field goals, how many two-point field goals did they make?

OpenStudy (anonymous):

3T + 2W = Total Pts

OpenStudy (anonymous):

b) 3(7) + 2W = 67 solve for W

OpenStudy (anonymous):

3(7) + 2W = 67 21 + 2W = 67 -21 + 21 + 2W = 67 + -21 2W = 46 divide each side by 2 W = 23 = # of 2pt field goals

OpenStudy (anonymous):

Thanks!